[ the hypotenuse of a triangle is 14 in. if the base of the triangle is 2 in determine the length of the remaining side ]
Answer by TimothyLamb(4379)

h = sqrt( b^2 + L^2 )
14 = sqrt( 2^2 + L^2 )
14^2 = 4 + L^2
L^2 = 196 - 4
L^2 = 192
---
L = 13.856406 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php